OpenStudy (anonymous):

Find the exact value of the six trig functions of the given angle. 300 degrees.

OpenStudy (anonymous):

sin300 = -sqrt(3)/2 cos300 =1/2 tan300 = -sqrt3

OpenStudy (anonymous):

cot300degrees? sec300degrees? csc300degrees?

OpenStudy (anonymous):

find the corresponding reciprocals